OpenStudy (anonymous):

help please 3x + 2 = 5x – 6 show steps please

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

first isolate the variables (x) in side and the constants (numbers) in the other side of the equal sign

OpenStudy (anonymous):

um

OpenStudy (anonymous):

is it clear ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what do you got ?

OpenStudy (anonymous):

just trying to under stand how to do that

OpenStudy (anonymous):

i know the answer is x=4

OpenStudy (anonymous):

i just forgot the steps

OpenStudy (anonymous):

3x + 2 = 5x – 6 move all the x's to one side by subtraction 3x + 2 = 5x – 6 -3x -3x which gives you 2 = 2x - 6 then isolate the x by adding 6 to both sides 2 = 2x - 6 +6 +6 then you have 8 = 2x divide both sides by 2 8/2 = 2x/2 which gives you 4 =x hope this helps!

OpenStudy (anonymous):

depending on the sign of the term if it was positive you subtract both sides by the value of the term and if it was negative you add the value of the term to the both sides

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

your welcome